# Hedgehog topological defects in 3D amorphous solids

### v1

@@ Abstract
In order to test the predictability of the plastic regions from pure static configurations, we obtain the soft spots using method (I) applied just before the plastic events. In Fig. 3(a)-(d) we show the spatial correlation between HTDs within non-affine displacement field (see SM for details) and the soft spots identified with the softness parameter $\phi_{i}$ (method I) for four plastic events at $\gamma=0.099,~{}0.120,~{}0.126$ and $0.154$ (see Fig. S1(c) for associated stress drop). From there, it is evident that both positive and negative HTDs with displacement field display a clear correlation with the soft spots that is rather localized around $r\approx 1$. Once again, no marked distinction between negative and positive defects is observed.
[…]
@@
In order to confirm this result, in the bottom panels of Fig.4, we show the spatial correlation function between monopoles in the displacement field and soft spots (red data) and between anti-monopoles in the displacement field and soft spots (blue data). For all the three values of strain considered, we observe a very strong spatial correlation between the HTDs in the displacement field and the soft spots, with a pronounced peak appearing at $r\approx 1$. Interestingly, the distance at which this correlation emerges is compatible with the low-frequency results presented in Fig. 2, where the HTDs have been identified in the eigenvector field, and the results in Fig. 3 as well. This confirms our initial speculation that, around plastic events, the topology of the low-frequency eigenvectors, and even its spatial characteristics, are compatible, and in fact very similar, to those observed in the displacement vector.
@@
Finally, we notice that there is no apparent difference between negative ($Q=-1$) and positive ($Q=+1$) HTDs, or at least in their correlation with plasticity. This is independent of the methods employed to define the HTDs and also the soft spots, and therefore rather universal. Interestingly, this represents a dissimilarity with respect to the analysis using the eigenvector field in 2D (or in the 2D-slicing method for 3D systems), where negative defects seem to have a privileged role Wu et al. (2023); Desmarchelier et al. (2024); Bera et al. (2024). In addition, we notice that the results shown in Fig. 4, using the non-affine displacement field, display a rather long-range structure in which spatial correlations survive up to large distances, $r\approx 5$. This very long correlation length might depend on several parameters and deserves further investigation. Moreover, the spatial structure of the HTDs and the spatial decay of their correlation with soft spots might provide useful insights to achieve a microscopic definition of the elastic screening length proposed in other macroscopic approaches to amorphous solids Lemaître et al. (2021).
@@
Figure 4: (a)-(c) The location of HTDs with charges $Q=+1$ and $Q=-1$ in the displacement vector field at different $\gamma$ are shown in red and colorblueblue, respectively. The yellow regions are the plastic soft spots, identified using the softness parameter, Eq. (6). (d)-(f) The spatial correlation between between monopoles (in the displacement field) and soft spots (red data) and anti-monopoles (in the displacement field) and soft spots (blue data) for the same values of the strain $\gamma$. For this analysis, soft spots were identified from the top $10\%$ of highest $|\vec{u}_{\text{NA}}|$ values.
@@
Conclusions – In this work, we proposed to use point-like hedgehog topological defects (HDTs) to characterize the plasticity of 3D amorphous solids and to locate their soft spots using well-defined mathematical concepts. First, borrowing from the physics of liquid crystals Kleman and Lavrentovich (2003, 2006); Hölbl et al. (2023) and Heisenberg magnets Fumeron and Berche (2023); Berg and Lüscher (1989); Holm and Janke (1994); Berg and Lüscher (1981), we have outlined a concrete and practical procedure to identify HDTs in 3D glasses – a method that is suitable for experimental discrete data sets as well.
@@
Finally, we considered HTDs in the 3D displacement vector field for a quasi-static shear deformation and for different large plastic events, including the yielding point. In this analysis, we have found that HTDs correlate very strongly with the extended plastic soft spots from which the plastic instability originates, that are defined using two different methods based respectively on the softness parameter Smessaert and Rottler (2014b) and the amplitude of the non-affine displacement field. This result proves that the location of HTDS, and their yet undisclosed dynamics, might play a pivotal role to achieve a microscopic understanding of plasticity and yielding in 3D glasses.
@@
Interestingly, the analysis of topological defects in 3D amorphous solids revealed that the correlation with the plastic soft spots is independent of the charge of the defect, unlike previous findings in 2D systems Wu et al. (2023); Desmarchelier et al. (2024); Bera et al. (2024) and in the 2D-slicing of 3D systems Bera et al. (2024). At present, we do not have an explanation for this intriguing phenomenon that requires further study.
@@
Before concluding, we also notice that the correlation between plasticity and topological defects appears stronger when the latter are defined using the displacement field, rather than the low-frequency eigenvector fields. This is promising since, apart from few exceptions (e.g., Vaibhav et al. (2024)), the eigenvector field is not a quantity directly accessible via experiments, whereas the displacement vector is.
@@
On a general ground, our work proves that the idea of using geometric topology to describe, predict and rationalize plastic properties of amorphous solids can be successfully extended to three dimensional systems, such as real materials. Aside from the theoretical part of this development, this result provides a formidable occasion to test these ideas on realistic experimental amorphous solids for which the method we proposed can be directly applied. For example, three dimensional sheared granular systems Cao et al. (2018) constitute a possible candidate platform to test these ideas.
Acknowledgments – M.B. thanks W. Kob, Z. Wu, J. Zhang, Y. Wang, C. Jiang, Z. Zheng, H. Tong for useful discussions. We would like to thank A. Liu and T. Petersen for ongoing collaborations and endless discussions about this topic as well. M.B. acknowledges the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01) and the sponsorship from the Yangyang Development Fund. A.Z. gratefully acknowledges funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. A.Z. gratefully acknowledges the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program. A.Z. and A.B. gratefully acknowledge funding from US Army Research Office through Contract No. W911NF-22-2-0256.
Supplementary Material
In this Supplementary Material (SM), we provide further details about the numerical and analytical computations presented in the main text. Moreover, we show further analysis to corroborate our findings.
Model and numerical methods
@@
In this study, we implement the Kremer-Grest model Kremer and Grest (1990) to simulate a coarse-grained polymer system. This system comprises linear polymer chains, each consisting of 50 monomers, with alternating monomer masses of $m_{1}=1$ and $m_{2}=3$ Kriuchevskyi et al. (2020). The total number of monomers in our system was $N=10{,}000$.
@@
Figure S1: (a) Schematic representation of the polymer chains with alternating masses $m_{1}$ and $m_{2}$. The FENE bonds and the LJ interactions range $r_{c}$ from a monomer (A) are shown. (b) A typical simulation configuration is presented where monomer breads with masses $m_{1}$ and $m_{2}$ are indicated in red and blue, respectively and FENE bonds by green colors. (c) Stress versus strain curve. The yielding point is indicated by the dashed black line at $\gamma=0.099$.
@@
$u_{\text{LJ}}(r)=\begin{cases}V(r)-V(r_{c}),&\text{for}~{}r<r_{c}=2.5\sigma,\\ 0,&\text{otherwise.}\end{cases}$ (S1)
@@
$V(r)=4{\varepsilon}\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right],$ (S2)
@@
$u_{\text{FENE}}(r)=-\frac{Kr_{0}^{2}}{2}\ln\left[1-\left(\frac{r}{r_{0}}\right)^{2}\right].$ (S3)
@@
We set $K=30$ and $r_{0}=1.5\sigma$ in our simulations. In Fig. S1(a) we schematically show these interactions. The polymer chains were simulated within a cubic box with periodic boundary conditions applied in all three dimensions. The system was equilibrated using the LAMMPS simulation package Plimpton (1995). We show a typical configuration in Fig. S1(b). Following equilibration, we employed an athermal quasi-static (AQS) deformation protocol. This involved quenching the glass sample to absolute zero temperature and subsequently applying quasi-static shear with a strain increment of $\delta\gamma_{xz}=0.001$. The units of mass, length, and time of our simulation are taken as $m_{1}$, $\sigma$, and $\tau=\sqrt{m_{1}\sigma^{2}/{\varepsilon}}$, respectively. We set $m_{1}$, $\sigma$, and $\varepsilon$ to unity. Throughout our simulations, polymer chains were considered fully flexible, with no angle-bending terms included in the potential. The stress-strain curve in the shear plane is presented in Fig. S1(c).
Calculation of smoothed eigenvector field
@@
Given the Hessian matrix, we compute the eigenvector field ${\mathbf{e}}_{k}=(e^{x}_{k},e^{y}_{k},e^{z}_{k})$ for each eigenmode with eigenfrequency $\omega_{k}$ ($k=1,2,\ldots,3N$). This eigenvector field is interpolated onto a $40\times 40\times 40$ simple cubic lattice with grid spacing of $\sigma/2$ in all directions, superimposed on the 3D simulation box. The eigenvector field at each lattice site $\vec{r}=(l,m,n)$ is then obtained as Wu et al. (2023)
$e^{\mu}_{k}(\vec{r})=\frac{\sum_{d_{i}<R_{\text{cut}}}w(\vec{r}-\vec{r}_{i})e^{\mu}_{i}}{\sum_{d_{i}<R_{\text{cut}}}w(\vec{r}-\vec{r}_{i})},\quad\mu=x,y,z,$ (S4)
where $\vec{r}_{i}$ is the position of particle $i$, and $w(\vec{r}-\vec{r}_{i})$ is a Gaussian weight function given by $w(\vec{r}-\vec{r}_{i})=\exp(-|\vec{r}-\vec{r}_{i}|^{2}/r_{c}^{2})$ with $r_{c}=1$. The distance $d_{i}=|\vec{r}-\vec{r}_{i}|$ is between the lattice site and the particle $i$, and $R_{\text{cut}}=4\sigma$ is the cut-off distance for interpolation.
@@
Figure S2: Normalized eigenvector field for a portion of the simulation box exhibiting a monopole defect (red) and an anti-monopole one (blue). These data correspond to $\omega=3$.
Calculation of topological charge density
@@
In order to compute the topological charge density, we divide the entire 3D simulation box into cubical cells of side $\sigma/2$. Within each such cell we count number of HTDs $N^{+}$ and $N^{-}$ with $Q=+1$ and $Q=-1$, respectively, for $\omega_{k}<4$. The topological charge density inside each cubical cell is defined as
$\bar{C}=\frac{\sum_{\omega_{k}}(N^{+}-N^{-})/{\omega}_{k}^{2}}{\sum_{\omega_{k}}1/{\omega}_{k}^{2}},$ (S5)

### v2

@@ Abstract
In order to test the predictability of the plastic regions from pure static configurations, we obtain the soft spots using method (I) applied just before the plastic events. We investigate the spatial correlation between HTDs within non-affine displacement field and the soft spots identified with the softness parameter $\phi_{i}$ (method I) across various plastic events. We analyzed $10$ plastic events, for which the stress drops in the stress-strain curve are indicated in the Methods section and shown in Figure 6c. In Figure 4a, we present the average correlations between soft spots (Method I) and HTDs with topological charges $Q=+1$ and $Q=-1$. From there, it is evident that both positive and negative HTDs within the displacement field display a clear correlation with the soft spots that is rather localized around $r\approx 1$. Once again, no marked distinction between negative and positive defects is observed. We now distinguish the HTDs based on their geometric projections, classifying them as hyperbolic or radial, as discussed earlier. In Figure 4b, we present the correlations between soft spots and hyperbolic defects ($g_{sH}(r)$), as well as soft spots and radial defects ($g_{sR}(r)$). We observe that the correlation for hyperbolic defects is stronger than for radial defects at shorter distances. These findings further support the observations discussed in the previous sections.
We then move on to perform the same analysis using method (II), based on the non-affine displacement magnitude. We present a snapshot at $\gamma=0.099$, highlighting the locations of soft spots and HTDs with hyperbolic and radial geometries. In Figure 5a-c, we show 2D projections in the $xy$, $yz$, and $xz$ planes, respectively, along with a full 3D view in Figure 5d. The strong spatial correlation between the soft spots and HTDs is evident.
We compute the spatial correlation function between monopoles in the displacement field and soft spots (red data) and between anti-monopoles in the displacement field and soft spots (blue data). In Figure 5e, we present the averaged correlation functions between soft spots (method II) and HTDs with topological charges $Q=+1$ and $Q=-1$, averaged across 10 different plastic events, as shown in Figure 6c. Additionally, in Figure 5f, we show the averaged correlation functions between soft spots (method II) and HTDs classified by hyperbolic and radial nature.
@@
We observe a very strong spatial correlation between the HTDs in the displacement field and the soft spots, with a pronounced peak appearing at $r\approx 1$. Interestingly, the distance at which this correlation emerges is compatible with the low-frequency results presented in Figure 2, where the HTDs have been identified in the eigenvector field, and the results in Figure 3 as well. This confirms our initial speculation that, around plastic events, the topology of the low-frequency eigenvectors, and even its spatial characteristics, are compatible, and in fact very similar, to those observed in the displacement vector.
@@
Finally, we notice that there is no apparent difference between negative ($Q=-1$) and positive ($Q=+1$) HTDs, or at least in their correlation with plasticity. This is independent of the methods employed to define the HTDs and also the soft spots, and therefore rather universal. Interestingly, this represents a dissimilarity with respect to the analysis using the eigenvector field in 2D (or in the 2D-slicing method for 3D systems), where negative defects seem to have a privileged role [45, 50, 54]. When we classify the HTDs based on their geometric nature as hyperbolic or radial, we observe stronger correlations for hyperbolic defects. This trend holds consistently across all cases studied in this work.
In addition, we notice that the results shown in Figure 4, using the non-affine displacement field, display a rather long-range structure in which spatial correlations survive up to large distances, $r\approx 5$. This very long correlation length might depend on several parameters and deserves further investigation. Moreover, the spatial structure of the HTDs and the spatial decay of their correlation with soft spots might provide useful insights to achieve a microscopic definition of the elastic screening length proposed in other mesoscopic approaches to amorphous solids [49].
@@
Figure 5: Visualization and correlation between HTDs and non-affine regions from displacement field. a-c The location of HTDs with hyperbolic and radial nature in the displacement vector field at $\gamma=0.099$ are shown in two dimensional projections ($xy$, $yz$ and $xz$ planes) in magenta (hyperbolic) and green (radial), respectively. The yellow regions are the plastic soft spots, identified using the softness parameter, Eq. (7). d The 3D configuration with defects locations and soft spots at $\gamma=0.099$ are shown. e The average spatial correlation between monopoles (in the displacement field) and soft spots and anti-monopoles (in the displacement field) and soft spots. f The average spatial correlation between hyperbolic defects (in the displacement field) and soft spots and radial defects (in the displacement field) and soft spots are shown in magenta and green respectively. The average is calculated over several $\gamma$ values corresponding to sudden stress drops, and error bars are included.
Discussion
@@
In this work, we proposed to use point-like hedgehog topological defects (HDTs) to characterize the plasticity of 3D amorphous solids and to locate their soft spots using well-defined mathematical concepts. First, borrowing from the physics of liquid crystals [57, 58, 59] and Heisenberg magnets [60, 61, 62, 63], we have outlined a concrete and practical procedure to identify HDTs in 3D glasses – a method that is suitable for experimental discrete data sets as well.
@@
Finally, we considered HTDs in the 3D displacement vector field for a quasi-static shear deformation and for different large plastic events, including the yielding point. In this analysis, we have found that HTDs correlate very strongly with the extended plastic soft spots from which plastic instabilities originate. These soft spots are identified using two different methods: one based on the softness parameter [70] and the other on the amplitude of the non-affine displacement field. This result provides strong evidence that the location of HTDs, along with their yet undisclosed dynamics, may play a pivotal role in achieving a microscopic understanding of plasticity and yielding in 3D glasses.
@@
Interestingly, the analysis of topological defects in 3D amorphous solids revealed that the correlation with the plastic soft spots is independent of the charge of the defect, unlike previous findings in 2D systems [45, 50, 54] and in the 2D-slicing of 3D systems [54]. This results can be directly understood from the arbitrariness in the sign of the topological charge $Q$.
On the other hand, we find that in 3D, the geometrical properties of the point defects play a fundamental role in identifying the microscopic structures responsible for plasticity. In particular, we observe that hedgehog topological defects with hyperbolic nature are ultimately the responsible for plastic deformations in 3D amorphous solids. This aligns well with previous findings in 2D [45], where the corresponding carriers are the ones with negative winding number (anti-vortices). The connection between these two objects is that they both share a saddle-like geometrical structure that is more prone to plastic instabilities and hence strongly correlates with the plastic spots.
@@
Before concluding, we also notice that the correlation between plasticity and topological defects appears stronger when the latter are defined using the displacement field, rather than the low-frequency eigenvector fields. This is promising since, apart from few exceptions (e.g., [51]), the eigenvector field is not a quantity directly accessible via experiments (with the exception of colloidal systems [51]), whereas the displacement vector is. Although there may be a connection between the structural properties of the material and these HTDs, none of the recent studies have explicitly reported it. See [75] for a recent attempt in this direction based on inversion-symmetry and [76] for a study in defective crystals.
@@
On a general ground, our work proves that the idea of using geometry and topology to describe, predict and rationalize plastic properties of amorphous solids can be successfully extended to three dimensional systems, such as real materials. Aside from the theoretical part of this development, this result provides a formidable occasion to test these ideas on realistic experimental amorphous solids for which the method we proposed can be directly applied. For example, three dimensional sheared granular [52] and colloidal [77] systems constitute possible candidate platforms to test these ideas.
Methods
Model
@@ Methods
In this study, we implement the Kremer-Grest model [66] to simulate a coarse-grained polymer system. This system comprises linear polymer chains, each consisting of 50 monomers, with alternating monomer masses of $m_{1}=1$ and $m_{2}=3$ [67]. The total number of monomers in our system was $N=10{,}000$.
@@
Figure 6: Simulation model and stress versus strain curve. a Schematic representation of the polymer chains with alternating masses $m_{1}$ and $m_{2}$. The FENE bonds and the LJ interactions range $r_{c}$ from a monomer (A) are shown. b A typical simulation configuration is presented where monomer breads with masses $m_{1}$ and $m_{2}$ are indicated in red and blue, respectively and FENE bonds by green colors. c Stress ($\sigma$) versus strain ($\gamma$) curve. Several stress drops, indicated as plastic events, are marked by black dashed lines. The yielding point ($\gamma=0.099$) and other plastic events at $\gamma=0.012$, $0.126$, $0.154$, $0.162$, $0.175$, $0.204$, $0.207$, $0.214$, and $0.238$ are analyzed separately, and the averaged correlation results are presented in Figure 4a-b and Figure 5e-f in the main text.
@@
$u_{\text{LJ}}(r)=\begin{cases}V(r)-V(r_{c}),&\text{for}\leavevmode\nobreak\ r<r_{c}=2.5\sigma,\\ 0,&\text{otherwise.}\end{cases}$ (9)
@@
$V(r)=4{\varepsilon}\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right],$ (10)
@@
$u_{\text{FENE}}(r)=-\frac{Kr_{0}^{2}}{2}\ln\left[1-\left(\frac{r}{r_{0}}\right)^{2}\right].$ (11)
@@
We set $K=30$ and $r_{0}=1.5\sigma$ in our simulations. In Figure 6a we schematically show these interactions. The polymer chains were simulated within a cubic box with periodic boundary conditions applied in all three dimensions. The system was equilibrated using the LAMMPS simulation package [78]. We show a typical configuration in Figure 6b. Following equilibration, we employed an athermal quasi-static (AQS) deformation protocol. This involved quenching the glass sample to absolute zero temperature and subsequently applying quasi-static shear with a strain increment of $\delta\gamma_{xz}=0.001$. The units of mass, length, and time of our simulation are taken as $m_{1}$, $\sigma$, and $\tau=\sqrt{m_{1}\sigma^{2}/{\varepsilon}}$, respectively. We set $m_{1}$, $\sigma$, and $\varepsilon$ to unity. Throughout our simulations, polymer chains were considered fully flexible, with no angle-bending terms included in the potential. The stress-strain curve in the shear plane is presented in Figure 6c.
Numerical methods and data analysis
@@
We diagonalize the dynamical matrix as described in Eq. 6 to obtain the eigenvalues and eigenvectors. For each eigenvector, represented as a $3N\times 1$ column matrix corresponding to an eigenfrequency $\omega_{k}$ ($k=1,2,\ldots,3N$), we assign the eigenvector field ${\mathbf{e}}_{i}=(e^{x}_{i},e^{y}_{i},e^{z}_{i})$ to each particle position $\vec{r}_{i}$ ($i=1,2,\ldots,N$). This eigenvector field is interpolated onto a $40\times 40\times 40$ simple cubic lattice with grid spacing of $\sigma/2$ in all directions, superimposed on the 3D simulation box. For each eigenfrequency, the eigenvector field at each lattice site $\vec{r}=(l,m,n)$ is then obtained as [45]
$e^{\mu}(\vec{r})=\frac{\sum_{d_{i}<R_{\text{cut}}}w(\vec{r}-\vec{r}_{i})e^{\mu}_{i}}{\sum_{d_{i}<R_{\text{cut}}}w(\vec{r}-\vec{r}_{i})},\quad\mu=x,y,z,$ (12)
where $\vec{r}_{i}$ is the position of particle $i$, and $w(\vec{r}-\vec{r}_{i})$ is a Gaussian weight function given by $w(\vec{r}-\vec{r}_{i})=\exp(-|\vec{r}-\vec{r}_{i}|^{2}/r_{c}^{2})$ with $r_{c}=1$. The distance $d_{i}=|\vec{r}-\vec{r}_{i}|$ is the separation between the lattice site and particle $i$, and $R_{\text{cut}}=4\sigma$ is the cutoff distance used for interpolation. We employ this cutoff radius, $R_{\text{cut}}$, to accelerate the simulation. We tested the robustness of the interpolated field by considering larger cutoff distances and the full long-range limit. The eigenvector field at each lattice grid point is then normalized and used to identify the hedgehog topological defects (HTDs). In Figure 7, we show a snapshot of a 3D eigenvector field presenting two clear topological defects with opposite charge at eigenfrequency $\omega=3$.
@@
Figure 7: Normalized eigenvector field for a portion of the simulation box exhibiting a monopole defect (red) and an anti-monopole one (blue). These data correspond to $\omega=3$.
@@
In order to compute the topological charge density, we divide the entire 3D simulation box into cubic cells of side $\sigma/2$. Within each such cell we count number of HTDs $N^{+}$ and $N^{-}$ with $Q=+1$ and $Q=-1$, respectively, for $\omega_{k}<4$. The topological charge density inside each cubic cell is defined as
$\bar{C}=\frac{\sum_{\omega_{k}}(N^{+}-N^{-})/{\omega}_{k}^{2}}{\sum_{\omega_{k}}1/{\omega}_{k}^{2}},$ (13)
@@
We derive the displacement field as [44, 54] $\vec{u}_{i}={\vec{r_{i}}}(\gamma)-{\vec{r}_{i}}(\gamma-\delta\gamma)$, where $\vec{r}_{i}(\gamma)$ and $\vec{r}_{i}(\gamma-\delta\gamma)$ represents the locations of the particle $i$ at strains $\gamma$ and $\gamma-\delta\gamma$, respectively. The affine component for the simple shear scenario is calculated as $\vec{u}_{A}=z\delta\gamma\hat{x}$. Finally we obtain the non-affine displacement field by subtracting the affine component from the total displacement, $\vec{u}_{\text{NA}}=\vec{u}_{i}-\vec{u}_{A}$ [44, 54]. To obtain HTDs within the non-affine displacement field, we interpolated the discrete non-affine displacement field $\vec{u}_{\text{NA}}$ onto a simple cubic lattice grid with a side length $\sigma/2$ using the standard Python package griddata with linear interpolation.
Data availability
The datasets generated and analysed during the current study are available upon reasonable request by contacting the corresponding authors.
Code availability
The codes that support the findings of this study is available upon reasonable request by contacting the corresponding authors.
Acknowledgements
M.B. thanks W. Kob, Z. Wu, J. Zhang, Y. Wang, C. Jiang, Z. Zheng, H. Tong for useful discussions. We would like to thank A. Liu and T. Petersen for ongoing collaborations and endless discussions about this topic as well. M.B. acknowledges the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01) and the sponsorship from the Yangyang Development Fund. A.Z. gratefully acknowledges funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. A.Z. gratefully acknowledges the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program. A.Z. and A.B. gratefully acknowledge funding from US Army Research Office through Contract No. W911NF-22-2-0256.
Author contributions
A. B. performed the numerical computations and data analysis. M. B. and A. B. wrote the manuscript with the help of A. Z. All Authors contributed to the physical interpretation of the results.
Competing interests
The authors declare no competing interests.
Supplementary Material